### v5

# Antipodality of Spherical Designs with Odd Harmonic Indices

Abstract.
We determine the smallest size of a non-antipodal spherical design with harmonic indices $\{1,3,\dots,2m-1\}$ to be $2m+1$, where $m$ is a positive integer. This is achieved by proving an analogous result for interval designs.
1991 Mathematics Subject Classification:
65D32,05E99
1. Introduction
Spherical designs appear in approximation theory, frame theory and combinatorics (see [1, 6, 18]). A more general concept of spherical designs was introduced by [5]. A finite subset $X$ of a unit sphere $\mathbb{S}^{d-1}$ in $\mathbb{R}^{d}$ is defined to be a spherical design of harmonic index $t$ if the sum of the values on $X$ vanishes for every harmonic homogeneous polynomial of degree $t$. If $T$ is a set of positive integers and $X$ is a spherical design of harmonic index $t$ for every $t\in T$, we simply say that $X$ is a spherical $T$-design. Then the ordinary spherical $t$-design is nothing but a spherical $\{1,2,\dots,t\}$-design. For certain finite sets $T$ of even integers, spherical $T$-designs with additional conditions on size or the number of distances have been considered (see, e.g., [7, 9, 14, 19]).
If $X$ consists of a pair of antipodal points on $\mathbb{S}^{d-1}$, then $X$ is a spherical design of harmonic index $t$ for every odd positive integer $t$. Conversely, it is not difficult to show that if $X\subset\mathbb{S}^{d-1}$ is a spherical design of harmonic index $t$ for all odd positive integers $t$, then $X$ is necessarily a union of pairs of antipodal points on $\mathbb{S}^{d-1}$. For a brief proof of this fact, suppose that there exists a point $x\in X$ with $x\notin X$. Since $|\langle x,y\rangle|<1$ for all $y\in X\setminus\{x\}$,
$\sum_{y\in X}\langle x,y\rangle^{2k-1}\to 1$
as $k\to\infty$. This contradicts the assumption that $X$ is a spherical design of harmonic index 2k12k-1 for all positive integers $k$.
Since a regular $(2m+1)$-gon in $\mathbb{S}^{1}$ is known to be a spherical $2m$-design (see [8]), it is also a spherical $\{1,3,\dots,2m-1\}$-design. Thus, if $T$ is a finite set of positive integers with $2m-1=\max T$, then there exists a $(2m+1)$-point spherical $T$-design which is not a union of pairs of antipodal points.
Our main result is to show that $2m+1$ is the smallest size of a finite set $X$ which can be a non-antipodal spherical $\{1,3,\dots,2m-1\}$-design. This is achieved by proving an analogous result for interval designs.
The paper is organized as follows. In Section 2, we recall Newton’s identities that relate power sums and elementary symmetric polynomials. In Section 3, we introduce an analogue of harmonic indices for interval designs and prove a lower bound for the size of a non-symmetric design. By regarding this result as a set version, we also prove its function version (i.e., a design with weights). In Section 4, we prove our main result on a lower bound for the size of a non-antipodal spherical design with odd harmonic indices. Finally, in Section 5, we prove the optimality of our results in Sections 3 and 4 by constructing non-antipodal designs with the smallest possible sizes.
2. Preliminaries
For each non-negative integer $k$, we denote by $p_{k}(x_{1},\dots,x_{n})$ the $k$-th power sum and by $e_{k}(x_{1},\dots,x_{n})$ the $k$-th elementary symmetric polynomial in the variables $x_{1},\dots,x_{n}$. The following fact, widely known as Newton’s identity, gives relations between elementary symmetric polynomials and power sums.
Lemma 2.1.
  1. (i)
    For all integers $n\geq k\geq 1$, we have
    $k\,e_{k}(x_{1},\ldots,x_{n})=\sum_{i=1}^{k}(-1)^{i-1}e_{k-i}(x_{1},\ldots,x_{n})\,p_{i}(x_{1},\ldots,x_{n}).$
  2. (ii)
    For all integers $k\geq n\geq 1$, we have
    $0=\sum_{i=k-n}^{k}(-1)^{i-1}e_{k-i}(x_{1},\ldots,x_{n})\,p_{i}(x_{1},\ldots,x_{n}).$
Since we will be dealing with spherical designs with odd harmonic indices, we establish relationships between power sums of odd degrees and elementary symmetric polynomials of odd degrees.
Proposition $2.$
Let $m$ and $n$ be positive integers with $2m-1\leq n$, and let $x_{1},\ldots,x_{n}\in\mathbb{R}$. Then the following are equivalent:
  1. (i)
    $p_{2k-1}(x_{1},\ldots,x_{n})=0\quad\text{for all }k\in\{1,\dots,m\}.$
  2. (ii)
    $e_{2k-1}(x_{1},\ldots,x_{n})=0\quad\text{for all }k\in\{1,\dots,m\}.$
Proof.
The equivalence of (i) and (ii) is clear when $k=1$. We prove (i) $\Rightarrow$ (ii) by induction on $m$. By Lemma 2.1 (i),
$e_{2k-1}(x_{1},\ldots,x_{n})=\frac{1}{2k-1}\sum_{i=1}^{2k-1}(-1)^{i-1}e_{2k-1-i}(x_{1},\ldots,x_{n})\,p_{i}(x_{1},\ldots,x_{n}).$
Note that in each term $e_{2k-1-i}\,p_{i}$ (for $1\leq i\leq 2k-1$), the indices have opposite parity. Hence, by (i) and the induction hypothesis, we obtain $e_{2k-1}(x_{1},\ldots,x_{n})=0$.
Next, we prove (ii) $\Rightarrow$ (i) by induction on $k$. By Lemma 2.1 (i),
$p_{2k-1}(x_{1},\ldots,x_{n})$ $=(2k-1)e_{2k-1}(x_{1},\ldots,x_{n})$
$\quad+\sum_{i=1}^{2k-2}(-1)^{i}e_{2k-1-i}(x_{1},\ldots,x_{n})\,p_{i}(x_{1},\ldots,x_{n}).$
Again, in each term $e_{2k-1-i}\,p_{i}$ (for $1\leq i\leq 2k-2$), the indices have different parity. Then by (ii) and the induction hypothesis, we have $p_{2k-1}(x_{1},\ldots,x_{n})=0$. ∎
3. Interval Designs
Let $(a,b)$ be an open interval in $\mathbb{R}$ and let $w(x)$ be a probability density function such that all moments $\int_{a}^{b}x^{s}\,w(x)dx$ are finite. Let $x_{1},\ldots,x_{n}\in(a,b)$. In numerical analysis, an integration formula of the type
$\frac{1}{n}\sum_{k=1}^{n}x_{k}^{s}=\int_{a}^{b}x^{s}\,w(x)dx,$
for $s=1,2,\ldots,t$, is called a strict Chebyshev-type quadrature of degree $t$ if all $x_{i}$ are distinct, and a Chebyshev-type quadrature of degree $t$ otherwise [10]. As exemplified by the Chebyshev-Gauss quadrature
$\frac{1}{n}\sum_{k=1}^{n}\Bigl(\cos\Bigl(\frac{2k\pi}{2n-1}\Bigr)\Bigr)^{s}=\int_{-1}^{1}x^{s}\,\frac{1}{\pi\sqrt{1-x^{2}}}\,dx,$
where $s$ ranges over $1,2,\dots,2n-1$, Chebyshev-type formulas for the ultraspherical measure $(1-x^{2})^{\lambda-1/2}\,dx/\int_{-1}^{1}(1-x^{2})^{\lambda-1/2}\,dx$, where $\lambda$ is a nonnegative integer, are a central object in the study of quadrature formulas. Meanwhile, the study of the configuration of points $x_{1},\dots,x_{n}$ has long captured the attention of researchers in combinatorics and related areas, where the term “interval design” is widely used for the point configuration; for example see Bajnok [2].
Definition 3.1.
A subset $X=\{x_{1},x_{2},\dots,x_{n}\}$ of the interval $[a,b]$ is called an interval $t$-design if
(1) $\frac{1}{n}\sum_{k=1}^{n}x_{k}^{s}=\frac{1}{b-a}\int_{a}^{b}x^{s}\,dx$
holds for all $s=1,2,\dots,t$. More generally, for $T\subset\mathbb{N}$, the set $X$ is called an interval design of index $T$ if (1) holds for every $s\in T$.
We will only consider interval designs of index $T_{m}$ or interval $T_{m}$-designs, for brevity on the interval [1,1][-1,1], where
(2) $T_{m}=\{2k-1\mid 1\leq k\leq m\}.$
In particular, the definition can be stated for any measure on [1,1][-1,1] by replacing the right-hand side of (1) with the corresponding moment integral. It may also be worth mentioning that (1) is equivalently written as $p_{s}(x_{1},\dots,x_{n})=0$ for an odd positive integer $s$.
We say that a finite subset $X\subset[-1,1]$ is symmetric if X=XX=-X. More generally, a multiset $X\subset[-1,1]$ is said to be symmetric if the multiplicity of $x$ in $X$ coincides with that of x-x for every $x\in X$.
Proposition 3.2.
Let $m$ and $n$ be positive integers with $n\leq 2m$. Suppose that real numbers $x_{1},\dots,x_{n}\in[-1,1]$ satisfy
(3) $p_{2k-1}(x_{1},\dots,x_{n})=0$
for all $k\in\{1,\dots,m\}$. Then the multiset $\{x_{1},\dots,x_{n}\}$ is symmetric. In particular, if a subset $X\subset[-1,1]$ is an interval $T_{m}$-design with $|X|=n$, then $X$ is symmetric.
Proof.
Without loss of generality, we may assume $0\notin X$, because removing zeros does not affect the conditions $\sum_{x\in X}x^{2k-1}=0$ for $1\leq k\leq m$. We first prove that (3) holds for every positive integer $k$ by induction on $k$. Since 3 is assumed for $k\leq m$, so suppose $k>m$. Then $2k-1\geq n$. Moreover, because there are only $n$ variables, $e_{j}(x_{1},\dots,x_{n})=0$ for all $j>n$; in particular, for $j>2m$. It follows from Lemma 2.1 (ii) that
$p_{2k-1}(x_{1},\dots,x_{n})$ $=\sum_{i=2k-1-n}^{2k-2}(-1)^{i}e_{2k-1-i}(x_{1},\dots,x_{n})\,p_{i}(x_{1},\dots,x_{n})$
$=\sum_{i=1}^{m}e_{2i-1}(x_{1},\dots,x_{n})\,p_{2(k-i)}(x_{1},\dots,x_{n})$
$\quad-\sum_{i=1}^{m}e_{2i}(x_{1},\dots,x_{n})\,p_{2(k-i)-1}(x_{1},\dots,x_{n}).$
By Proposition 2.2 and (3), we have $e_{2i-1}(x_{1},\dots,x_{n})=0$ for $1\leq i\leq m$. Thus, the first term is zero. By the inductive hypothesis, the second term is also zero. Therefore, we have shown that (3) holds for every positive integer $k$.
We now prove the assertion by induction on $n$. The cases $n=1,2$ are obvious. Let $Y=\{y_{1},\dots,y_{n^{\prime}}\}$ be the set of all distinct elements of $x_{1},\dots,x_{n}$, that is, $Y=\{x_{1},\dots,x_{n}\}$ and $|Y|=n^{\prime}\leq n$. Define $A\in\mathbb{R}^{n^{\prime}\times n^{\prime}}$ by
$A_{i,j}=y_{j}^{2i-1}\quad(1\leq i,j\leq n^{\prime}).$
Define a column vector $\boldsymbol{u}$ whose $j$-th entry $u_{j}$ is given by
$u_{j}=|\{i\mid 1\leq i\leq n,\;x_{i}=y_{j}\}|\quad(1\leq j\leq n^{\prime}).$
Then the $k$-th entry of $A\boldsymbol{u}$ is
$\sum_{j=1}^{n^{\prime}}y_{j}^{2k-1}u_{j}=p_{2k-1}(x_{1},\dots,x_{n})=0,$
since we have shown that (3) holds for every positive integer $k$. Thus $A\boldsymbol{u}=0$, and hence
$0=\det A=\prod_{k=1}^{n^{\prime}}y_{k}\prod_{1\leq i<j\leq n^{\prime}}(y_{j}^{2}-y_{i}^{2}).$
This implies that, either there exists $j$ such that $y_{j}=0$, or there exist $i,j$ with $i<j$ such that $y_{i}=-y_{j}$. In the latter case, we may assume without loss of generality $x_{n-1}=-x_{n}$. Since we have shown that (3) holds for every positive integer $k$, we obtain
$p_{2k-1}(x_{1},\dots,x_{n-2})=p_{2k-1}(x_{1},\dots,x_{n})=0$
for every positive integer $k$. By induction, the multiset $\{x_{1},\dots,x_{n-2}\}$ is symmetric, and so is $\{x_{1},\dots,x_{n}\}$. The former case can be proved in a similar manner. ∎
Remark $3.$
Although in this paper we restrict ourselves to interval designs for the normalized Lebesgue measure on [1,1][-1,1], the notion of an interval design naturally extends to an arbitrary measure on an interval. In this broader setting, a challenging problem is the construction of interval designs with rational points and weights for the ultraspherical measure (see Bannai et al. [4, p. 208, Problem 2]). It was Mishima et al. [13] who first observed Proposition 3.2 for $m\leq 3$ and thereby established, together with explicit constructions, a classification theorem of rational 55-designs with 66 points for the Chebyshev measure $(1-t^{2})^{-1/2}dt/\pi$.
In Proposition 3.2, the set $\{x_{1},\dots,x_{n}\}$ is regarded as a multiset. This is equivalent to consider a function $\lambda$ on $\{x_{1},\dots,x_{n}\}$ taking values in the set of positive integers. If we allow $\lambda$ to be a real-valued function, we can also conclude $f$ to be an even function under a stronger hypothesis.
Definition 3.4.
Let $m$ be a positive integer and let $\lambda\colon[a,b]\to\mathbb{R}$ be a function with finite support $X\subset[a,b]$. The function $\lambda$ is called a weighted interval design of index $T$ if
(4) $\sum_{x\in X}x^{s}\,\lambda(x)=\frac{1}{b-a}\int_{a}^{b}x^{s}\,dx$
holds for all $s\in T$. The nonzero values $\lambda(x)$ are called weights.
Similar to interval designs, we will only consider weighted interval designs of index $T_{m}$, or weighted interval $T_{m}$-design for brevity, on the interval [1,1][-1,1].
We say that a function $\lambda:[-1,1]\to\mathbb{R}$ with finite support is symmetric if $\lambda(x)=\lambda(-x)$ holds for all $x\in[-1,1]$. The same terminology is used in [10, Section 6] in the context of the theory of quadrature formulas.
Lemma 3.5.
The set of all weighted interval designs of index $T_{m}$ forms a vector space $D_{m}$ which contains every symmetric function $\lambda:[-1,1]\to\mathbb{R}$ with finite support.
Proof.
This follows immediately from the definition. ∎
Proposition 3.6.
Let $m$ and $n$ be positive integers with $n\leq m$. Let $\lambda\colon\mathbb{R}\to\mathbb{R}$ be a function with finite support $X\not\ni 0$ of size $|X|=n$. Suppose that $\lambda$ is a weighted interval $T_{m}$-design. Then $n$ is even and $\lambda$ is symmetric.
Proof.
We prove the assertion by induction on $n$. If $n=1$, then it is clear that there is no such weighted interval $T_{m}$-design for $m\geq 1$. Hence, we may suppose $n\geq 2$. Let $X=\{x_{1},\dots,x_{n}\}$, and define $A\in\mathbb{R}^{n\times n}$ by
$A_{i,j}=x_{j}^{2i-1}\quad(1\leq i,j\leq n).$
Define $v\in\mathbb{R}^{n}$ by
$v_{j}=\lambda(x_{j})\quad(1\leq j\leq n).$
Then for $1\leq i\leq n$,
$(Av)_{i}$ $=\sum_{j=1}^{n}x_{j}^{2i-1}\lambda(x_{j})$
$=\sum_{x\in\mathbb{R}}x^{2i-1}\lambda(x)$
$=0.$
This implies
$0=\det A=\prod_{k=1}^{n}x_{k}\prod_{1\leq i<j\leq n}(x_{j}^{2}-x_{i}^{2}).$
Thus, there exist $i,j$ with $1\leq i<j\leq n$ such that $x_{j}=-x_{i}$. We may assume without loss of generality $x_{n}=-x_{n-1}$.
Suppose $n=2$. Then $x_{2}=-x_{1}$, and hence $x_{1}(\lambda(x_{1})-\lambda(x_{2}))=0$ which implies $\lambda(x_{1})=\lambda(-x_{1})$. Since the support of $\lambda$ is $\{x_{1},x_{2}\}=\{x_{1},-x_{1}\}$, we conclude $\lambda(x)=\lambda(-x)$ for all $x\in\mathbb{R}$.
Now suppose $n\geq 3$. Define a function $\lambda^{\prime}:=\lambda-\omega$, where
[…]
Then $\omega$ is symmetric with finite support, and the support of $\lambda^{\prime}$ has size either n1n-1 or n2n-2. Since $n\leq m$, we have $\lambda\in D_{m}\subset D_{n-1}$. By Lemma 3.5, it follows that $\omega\in D_{n-1}$, and $\lambda^{\prime}=\lambda-\omega\in D_{n-1}$. Then, the inductive hypothesis implies $\lambda^{\prime}$ is symmetric. Since $\lambda=\lambda^{\prime}+\omega$, $\lambda$ is also symmetric. In particular, $n$ is even. ∎
4. Spherical Designs
For $a,b\in\mathbb{R}^{d}$, we denote by $\langle a,b\rangle$ the standard inner product.
Definition 4.1.
For a positive integer $t$, a finite nonempty subset $X\subset\mathbb{S}^{d-1}$ is called a spherical $t$-design if
$\sum_{x\in X}\langle x,a\rangle^{k}=\begin{cases}\displaystyle\frac{1\cdot 3\cdots(2k-1)\langle a,a\rangle^{\frac{k}{2}}}{(d+2)(d+4)\cdots(d+2k-2)}\,&\text{if }k\in\{2,4,\ldots,2\lfloor\frac{t}{2}\rfloor\},\\[5.69054pt] 0,&\text{if }k\in\{1,3,\ldots,2\lfloor\frac{t-1}{2}\rfloor+1\},\end{cases}$
for all $a\in\mathbb{R}^{d}$.
Among various definitions of spherical designs (See [3, Theorem 2.2]), the above definition, which originally goes back to Hilbert’s solution of Waring problem, can be found in Lyubich and Vaserstein [11], Reznick [15]. Venkov [17, Section 3] also shows that the above definition is equivalent to the standard one in Delsarte, Goethals and Seidel  [6]. In fact, denote by $|\mathbb{S}^{d-1}|$ the surface area $\int_{\mathbb{S}^{d-1}}1\,d\rho$ of the sphere $\mathbb{S}^{d-1}$. It follows, by the $O(d)$-invariance of the uniform measure $\rho$ on $\mathbb{S}^{d-1}$, that
$\frac{1}{|\mathbb{S}^{d-1}|}\int_{\mathbb{S}^{d-1}}\langle y,a\rangle^{2k}d\rho=\frac{1}{|\mathbb{S}^{d-1}|}\int_{\mathbb{S}^{d-1}}\langle y,\frac{a}{\|a\|}\rangle^{2k}\langle a,a\rangle^{k}d\rho$
$=\frac{\langle a,a\rangle^{k}}{|\mathbb{S}^{d-1}|}\int_{\mathbb{S}^{d-1}}y_{1}^{2k}d\rho=\frac{1\cdot 3\cdots(2k-1)}{d(d+2)\cdots(d+2k-2)}\langle a,a\rangle^{k},$
which provides an interpretation of the right-hand side of the equation in Definition 4.1 in an analytic manner. For a brief explanation on the relation between spherical designs and Hilbert’s work, we refer the reader to Seidel [16].
The following more general concept of spherical designs was introduced by [5], and in its full generality by [12]. See also [7, 9, 19].
Definition 4.2.
Let $Q_{d,t}(s)$ denote the Gegenbauer polynomial of degree $t$ on $\mathbb{S}^{d-1}$ (see [5, Lemma 2.1]). Let $T$ be a set of positive integers. A finite nonempty subset $X\subset\mathbb{S}^{d-1}$ is called a spherical design of harmonic index $T$ if
$\sum_{x,y\in X}Q_{d,t}(\langle x,y\rangle)=0\quad\text{for all }t\in T.$
We shall only consider spherical designs of harmonic index $T_{m}$ (or spherical $T_{m}$-designs), where $T_{m}$ is defined in (2). It is clear that any pair of antipodal points forms a spherical $T_{m}$-design for every positive integer $m$. Thus, if $X$ is a spherical $T_{m}$-design containing such a pair $\{a,-a\}$, then $X\setminus\{a,-a\}$ is also a spherical $T_{m}$-design.
Lemma 4.3.
Let $X$ be a finite subset of $\mathbb{S}^{d-1}$. Then $X$ is a spherical $T_{m}$-design if and only if $\{\langle x,a\rangle\}_{x\in X}$ is an interval $T_{m}$-design for all $a\in\mathbb{S}^{d-1}$.
Proof.
Considering the univariate polynomial ring $\mathbb{R}[s]$, note that the span of the monomials $s^{t}$ for $t\in T_{m}$ coincides with that of the Gegenbauer polynomials $Q_{d,t}(s)$ (since $Q_{d,t}(s)$ is odd of degree $t$ when $t\in T_{m}$). This equivalence establishes the claim. ∎
Theorem $4.$
Let $X$ be a spherical $T_{m}$-design consisting of $n$ points. If $n\leq 2m$, then $X$ is antipodal.
Proof.
Suppose $X$ is not antipodal. Then there exists some $a\in X$ such that $a\notin X$. The projection of $X$ onto the line determined by $a$ yields a multiset of $n$ real numbers that is non-symmetric. By Proposition 3.2, such a configuration forces $n>2m$, contradicting the assumption. ∎
5. Optimality
In this section, we show that the inequalities in the assumptions of Propositions 3.2, 3.6 and Theorem 4.4 cannot be weakened. First, we show that the inequality $n\leq 2m$ in Proposition 3.2 cannot be weakened. Recall that we denote by $p_{k}$ and $e_{k}$ the $k$-th power sum and $k$-th elementary symmetric polynomial, respectively with $n$ variables. For a finite set $A=\{a_{1},\dots,a_{n}\}$ with $n$ elements, we write $e_{k}(a_{1},\dots,a_{n})=e_{k}(A)$ and $p_{k}(a_{1},\dots,a_{n})=p_{k}(A)$.
Proposition 5.1.
Let $m$ be a positive integer. Then for every positive integer $n$ with $n>2m$, there exists an interval $T_{m}$-design $X\subset[-1,1]$ with $|X|=n$ and $X\neq-X$. In particular, $2m+1$ is the smallest size of a non-symmetric interval $T_{m}$-design.
Proof.
It suffices to prove the case where $n=2m+1$. Indeed, a set $X$ with $|X|=2m+1$ satisfying the conclusion must satisfy the stronger property $X\cap(-X)=\emptyset$, by Proposition 3.2. Thus, the case $n=2m+2$ follows by adding 0 to $X$. For $n>2m+2$, we can construct a desired $T_{m}$-design by adding $\left\lfloor{(n-2m)/2}\right\rfloor$ pairs of antipodal points in (1,1)(-1,1) to $X$ with $|X|=2m+1$ or $2m+2$.
Let
$A=\left\{\pm\frac{2k-1}{2m}\mid 1\leq k\leq m\right\}.$
Let $f(x)$ be the monic polynomial of degree $2m$ having AA as its set of roots. We can choose a small $\varepsilon>0$ in such a way that the polynomial $g(x)=f(x)+\varepsilon$ has $2m$ roots in $(-1+\frac{1}{2m},1-\frac{1}{2m})\setminus\{\frac{1}{2m}\}$. Let $A^{\prime}$ be the set of roots of $g(x)$. Then, by the relation between roots and coefficients,
$e_{k}(A)=e_{k}(A^{\prime})\quad(0\leq k\leq 2m-1).$
By Lemma 2.1 using induction on $k$, we find
(5) $p_{k}(A)=p_{k}(A^{\prime})\quad(0\leq k\leq 2m-1).$
Since AA and $(A-\tfrac{1}{2m})\cup\{1\}$ are both symmetric, $p_{\mathrm{odd}}(A)=0$, and
$0$ $=p_{2k+1}((A-\frac{1}{2m})\cup\{1\})$
$=1+\sum_{a\in A}(a-\frac{1}{2m})^{2k+1}$
$=1+\sum_{a\in A}\sum_{\ell=0}^{2k+1}\binom{2k+1}{\ell}a^{\ell}(-\frac{1}{2m})^{2k+1-\ell}$
$=1+\sum_{\ell=0}^{2k+1}\binom{2k+1}{\ell}(-\frac{1}{2m})^{2k+1-\ell}p_{\ell}(A)$
$=1+\sum_{\ell=0}^{k}\binom{2k+1}{2\ell}(-\frac{1}{2m})^{2(k-\ell)+1}p_{2\ell}(A)$
$=1-\sum_{\ell=0}^{k}\binom{2k+1}{2\ell}\frac{1}{(2m)^{2(k-\ell)+1}}p_{2\ell}(A).$
Thus, by (5), we obtain
$1-\sum_{\ell=0}^{k}\binom{2k+1}{2\ell}\frac{1}{(2m)^{2(k-\ell)+1}}p_{2\ell}(A^{\prime})=0\quad(0\leq k\leq m-1).$
Since $A^{\prime}$ is also symmetric, we can reverse the above calculation with AA replaced by $A^{\prime}$, and we conclude
$p_{2k+1}((A^{\prime}-\frac{1}{2m})\cup\{1\})=0\quad(0\leq k\leq m-1).$
This implies that the set $X=(A^{\prime}-\frac{1}{2m})\cup\{1\}$ is an interval $T_{m}$-design with $|X|=2m+1$. Since $1\in X$ and $1\notin X$, we have $X\neq(-X)$.
The second assertion then follows from Proposition 3.2. ∎
Next, we show that the inequalities in the assumptions of Proposition 3.6 cannot be weakened. The following two propositions show that $m+1$ is the smallest size of the support of a non-symmetric weighted $T_{m}$-design.
Proposition 5.2.
Let $n$ be a positive integer. Define the function $\lambda\colon[-1,1]\to\mathbb{R}$ by
$\lambda(x)=\begin{cases}2,&\text{if }x=\cos\Bigl(\dfrac{2j\pi}{2{n}+1}\Bigr)\text{ for }1\leq j\leq{n},\\[2.84526pt] 1,&\text{if }x=1,\\[2.84526pt] 0,&\text{otherwise.}\end{cases}$
Then, $\lambda$ is a weighted interval $T_{{n}-1}$-design.
Proof.
For $1\leq l\leq n$,
$\sum_{x\in\operatorname{supp}(f)}x^{2l-1}\lambda(x)$ $=\sum_{j=1}^{n}2\cos^{2l-1}\Bigl(\frac{2j\pi}{2n+1}\Bigr)+1$
$=\sum_{j=0}^{2n}\cos^{2l-1}\Bigl(\frac{2j\pi}{2n+1}\Bigr)=0,$
since the vertices of the regular $(2n+1)$-gon form a spherical $2n$-design. ∎
In this way, we always obtain a function $\lambda$ with support size $n$ that is a weighted interval $T_{n-1}$-design. This shows that the inequality $n\leq m$ in the assumption of Proposition 3.6 cannot be relaxed.
Using a combinatorial lemma, we can also construct a non-symmetric weighted $T_{n-1}$-design on [1,1][-1,1] supported on $n$ rational nodes with rational weights. In its proof, we denote by $\delta_{ij}$ the Kronecker’s delta symbol.
Lemma 5.3.
For positive integers $n$ and $s$ with $0\leq s<n$, we have
$\sum_{j=1}^{n}(-1)^{j}j^{2s}\binom{2n}{n-j}=\begin{cases}-\binom{2n-1}{n-1}&\text{if $s=0$,}\\ 0&\text{if $1\leq s<n$.}\end{cases}$
Proof.
We prove the assertion by induction on $s$.
Suppose $s=0$, then for all $n>0$ we have:
$\sum_{j=1}^{n}(-1)^{j}\binom{2n}{n-j}$ $=\frac{1}{2}\sum_{j=1}^{n}(-1)^{j}\left(\binom{2n}{n-j}+\binom{2n}{n+j}\right)$
$=\frac{(-1)^{n}}{2}\sum_{j=0}^{2n}(-1)^{j}\binom{2n}{j}-\frac{1}{2}\binom{2n}{n}$
$=-\binom{2n-1}{n-1}$.
Now let $s\geq 1$, and assume that the assertion holds for s1s-1 and for all n>s1n>s-1. Since
$j^{2}\binom{2n}{n-j}$ $=n^{2}\binom{2n}{n-j}-2n(2n-1)\binom{2n-2}{n-1-j}\quad(1\leq j\leq n-1),$
we have for $n>s$,
$\sum_{j=1}^{n}(-1)^{j}j^{2s}\binom{2n}{n-j}$
$=(-1)^{n}n^{2s}+\sum_{j=1}^{n-1}(-1)^{j}j^{2s}\binom{2n}{n-j}$
$=(-1)^{n}n^{2s}+\sum_{j=1}^{n-1}(-1)^{j}j^{2s-2}n^{2}\binom{2n}{n-j}$
$\quad-2n(2n-1)\sum_{j=1}^{n-1}(-1)^{j}j^{2s-2}\binom{2n-2}{n-1-j}$
$=n^{2}\sum_{j=1}^{n}(-1)^{j}j^{2(s-1)}\binom{2n}{n-j}-2n(2n-1)\sum_{j=1}^{n-1}(-1)^{j}j^{2(s-1)}\binom{2(n-1)}{n-1-j}$.
By the inductive hypothesis, we have
$\sum_{j=1}^{n}(-1)^{j}j^{2s}\binom{2n}{n-j}$ $=-n^{2}\delta_{s,1}\binom{2n-1}{n-1}+2n(2n-1)\delta_{s,1}\binom{2n-3}{n-2}=0.$
Proposition 5.4.
Let $n$ be a positive integer. Define $\lambda\colon[-1,1]\to\mathbb{R}$ by
$\lambda(x)=\begin{cases}2j\binom{2n}{n-2j}&\text{if $x=\frac{2j}{n+1}$ for $1\leq j\leq\left\lfloor n/2\right\rfloor$,}\\ (2j-1)\binom{2n}{n-2j+1}&\text{if $x=-\frac{2j-1}{n+1}$ for $1\leq j\leq\left\lceil n/2\right\rceil$,}\\ 0&\text{otherwise.}\end{cases}$
Then $\lambda$ is a weighted interval $T_{n-1}$-design.
Proof.
For $1\leq k\leq n-1$, we have
$\sum_{x\in\mathbb{R}}x^{2k-1}\lambda(x)$
$=\frac{1}{(n+1)^{2k-1}}\left(\sum_{j=1}^{\left\lfloor n/2\right\rfloor}(2j)^{2k}\binom{2n}{n-2j}-\sum_{j=1}^{\left\lceil n/2\right\rceil}(2j-1)^{2k}\binom{2n}{n-2j+1}\right)$
$=\frac{1}{(n+1)^{2k-1}}\sum_{j=1}^{n}(-1)^{j}j^{2k}\binom{2n}{n-j}$
$=0,$
by Lemma 5.3. ∎
Finally, we show that the inequality $n\leq 2m$ in Theorem 4.4 cannot be weakened. We need some preparations.
Lemma $5.$
Let $m$ be a positive integer. If $X$ is a spherical $T_{m}$-design on $\mathbb{S}^{d-1}$, then $X$ is also a spherical $T_{m}$-design on $\mathbb{S}^{d^{\prime}-1}$ for $d^{\prime}>d$.
Proof.
This is immediate from Lemma 4.3, since the inner product $\langle x,a\rangle$ for vectors $x,a\in\mathbb{R}^{d}$ remains the same after regarding them as vectors in $\mathbb{R}^{d^{\prime}}$. ∎
Proposition 5.6.
If $n>2m$ and $n$ is odd, then there exists a non-antipodal spherical $T_{m}$-design with $n$ points. In particular, $2m+1$ is the smallest size of a non-antipodal spherical $T_{m}$-design.
Proof.
Let $X$ be a regular $(2m+1)$-gon in $\mathbb{S}^{1}$. Then $X$ is a spherical $2m$-design (see [8]), and hence a spherical $T_{m}$-design. For $n>2m+1$ and $n$ is odd, pick n2m1n-2m-1 pairs of antipodal points in $\mathbb{S}^{1}$ disjoint from $X$. Adding these pairs of points to $X$ gives a desired $T_{m}$-design with $n$ points. Since $|X|$ is odd, $X$ is not antipodal.
Furthermore, by Lemma 5.5, a spherical $T_{m}$-design in $\mathbb{S}^{1}$, when embedded in $\mathbb{S}^{d-1}$, is also a spherical $T_{m}$-design.
The second assertion then follows from Theorem $4.$ ∎
In the proof of Proposition 5.6, we used the natural embedding of $\mathbb{S}^{1}$ into $\mathbb{S}^{d-1}$. By [5, Theorem 1.1], a more elaborate embedding of a regular $(2m+1)$-gon into $\mathbb{S}^{d-1}$ can be used to yield a spherical design with harmonic index $2m$.
Although Proposition 5.6 shows that the inequality $n\leq 2m$ in the assumption of Theorem 4.4 cannot be relaxed, it does not assert the existence of a non-antipodal spherical $T_{m}$-design with $n$ points for even $n>2m$. We show that Proposition 5.6 cannot be strengthened to include the case where $n$ is even, by proving the non-existence of a non-antipodal $T_{2}$-design with 66 points on $\mathbb{S}^{1}$. Note that Proposition 5.6 does imply the existence of a non-antipodal $T_{m}$-design with $n\geq 4m+2$ points as a union of two disjoint $(2m+1)$-gons.
Proposition 5.7.
There is no non-antipodal $T_{2}$-design on $\mathbb{S}^{1}$ of the cardinality 66.
Proof.
Let $X\subset\mathbb{S}^{1}$ be a spherical $T_{2}$-design with $|X|=6$ points. We will show that $X$ is antipodal. Suppose not. Then $X\cap(-X)=\emptyset$ by Theorem $4.$ In particular, we may assume by rotating that $(\pm x_{1},y_{1})\in X$ for some $x_{1}$ with $0<x_{1}<1$, $0<y_{1}$. The projection $\bar{X}$ of $X$ onto the $x$-coordinate gives a weighted interval $T_{2}$-design, and by removing $\pm x_{1}$ with multiplicity 11, we obtain a weighted interval $T_{2}$-design $X^{\prime}$ consisting of 44 points. By Proposition 3.2, these 44 points form a symmetric multiset $X^{\prime}$. Suppose first that 44 points of $X^{\prime}$ are not distinct. If there are only 33 distinct points, then one of the points is 0 with multiplicity 22, which implies $(0,\pm 1)\in X$. This contradicts the fact that $X\cap(-X)=\emptyset$. If there are only two distinct points, then we have $X^{\prime}=\{\pm x_{2}\}$, which also contradicts the fact that $X\cap(-X)=\emptyset$. We next suppose that $X^{\prime}$ consists of 44 distinct points $\{\pm x_{2},\pm x_{3}\}$. Then we may assume
$X=\{(\pm x_{1},y_{1}),(x_{2},y_{2}),(-x_{2},\epsilon_{2}y_{2}),(x_{3},y_{3}),(-x_{3},\epsilon_{3}y_{3})\},$
where $\epsilon_{2},\epsilon_{3}\in\{\pm 1\}$, $x_{2}^{2}+y_{2}^{2}=x_{3}^{2}+y_{3}^{2}=1$, $y_{2}>y_{3}>0$. Since $X\cap(-X)=\emptyset$, we have $\epsilon_{2}=1$ and similaly $\epsilon_{3}=1$. But then $y_{1}+y_{2}+y_{3}=0$ and $y_{1}^{3}+y_{2}^{3}+y_{3}^{3}=0$. This cannot occur, since $y_{1}y_{2}y_{3}\neq 0$. ∎
Acknowledgements
The authors would like to thank Eiichi Bannai, Hiroshi Nozaki, Kohji Tasaka, and Masatake Hirao for helpful discussions. The third author is supported in part by the Early Support Program for Grant-in-Aid for Scientific Research (B) of Kobe University.